# Blind Spot Navigation in LLM Reasoning with Thought Space Explorer

### v2

@@ Appendix B Detailed Information of Case Study
B.1 Case Analysis
As our approach significantly enhances the reasoning capabilities of thought structures while maintaining minimal computational cost, it effectively addresses gaps within existing structures. Figure 5 provides a specific example of this enhancement by comparing the original thought chains with the new chains generated by our method. In the provided case study, we see the reasoning prompt “Explain why it is important for children to learn mathematics”. The original thought chains (Thought Chain 1 and Thought Chain 2) are linear and somewhat limited in scope, focusing and repeating on foundational aspects like “Critical Thinking”, “Academic Success” and “Career Opportunities”. However, the new chain generated by our method explores additional dimensions, such as ‘‘Enhancement of Cognitive and Memory Skills” and “Preparation for the Digital Age”. These new perspectives enrich the argument and provide a more comprehensive understanding of the importance of mathematics. For the detailed information of this case, refer to Appendix B.

### v3

@@ Appendix B Detailed Information of Case Study
Case Analysis
As our approach significantly enhances the reasoning capabilities of thought structures while maintaining minimal computational cost, it effectively addresses gaps within existing structures. Figure 6 and Table 4 provides a specific example of this enhancement by comparing the original thought chains with the new chains generated by our method. In the provided case study, we see the reasoning prompt “Explain why it is important for children to learn mathematics”. The original thought chains (Thought Chain 1 and Thought Chain 2) are linear and somewhat limited in scope, focusing and repeating on foundational aspects like “Critical Thinking”, “Academic Success” and “Career Opportunities”. However, the new chain generated by our method explores additional dimensions, such as ‘‘Enhancement of Cognitive and Memory Skills” and “Preparation for the Digital Age”. These new perspectives enrich the argument and provide a more comprehensive understanding of the importance of mathematics.